P1=21,P2=43,P3=41,P4=81
Each solves independently
To find: P(at least one solves)
=1−P(none solves)
Plan: 1. Pˉi=1−Pi
2. P(none)=Pˉ1Pˉ2Pˉ3Pˉ4 (independent)
3. P(≥1)=1−P(none)
Pˉi=1−Pi
Pˉ1=21,Pˉ2=41,Pˉ3=43,Pˉ4=87
P(none solves)=Pˉ1×Pˉ2×Pˉ3×Pˉ4
=21×41×43×87
=2×4×4×81×1×3×7
=25621
P(at least one)=1−25621
=256256−21=256235
P(at least one solves)=256235
Correct option: (1)